What is the speed of an apple dropped from a tree after 1.5 second? What distance will it cover during this time? Take g=10m/s2.

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago
DEAR STUDENT

Let the velocity by which the Apple is moving me v ms.

 

Initial Velocity of the speed(u) = 0

Time taken by the Apple to fall = 1.5 s.

 

Using the Galileo's First Equation of the motion,

v - u = at

∴ v - 0 = 10 × 1.5

⇒ v = 15 m/s.


Hence, the velocity of the Apple is 15 m/s.

THANKS